# Mixture of Structural-and-Textual Retrieval over Text-rich Graph Knowledge Bases

### v3

@@ Abstract
Text-rich Graph Knowledge Bases (TG-KBs) have become increasingly crucial for answering queries by providing textual and structural knowledge. However, current retrieval methods often retrieve these two types of knowledge in isolation without considering their mutual reinforcement and some hybrid methods even bypass structural retrieval entirely after neighboring aggregation. To fill in this gap, we propose a Mixture of Structural-and-Textual Retrieval (MoR) to retrieve these two types of knowledge via a Planning-Reasoning-Organizing framework. In the Planning stage, MoR generates textual planning graphs delineating the logic for answering queries. Following planning graphs, in the Reasoning stage, MoR interweaves structural traversal and textual matching to obtain candidates from TG-KBs. In the Organizing stage, MoR further reranks fetched candidates based on their structural trajectory. Extensive experiments demonstrate the superiority of MoR in harmonizing structural and textual retrieval with insights, including uneven retrieving performance across different query logics and the benefits of integrating structural trajectories for candidate reranking. Our code is available at https://github.com/Yoega/MoR.
@@ 1 Introduction
Since supporting knowledge in TG-KBs typically exhibits in both the textual and structural formats Jin et al. (2024b); Kolomiyets and Moens (2011), retrieval methods should be tailored to both formats effectively as shown in Figure 1(a). Textual retrieval methods retrieve textual knowledge such as indexed documents based on its similarity to the given query and can be broadly categorized into lexical methods (e.g., BM25) and semantic methods (e.g., Contriever) Karpukhin et al. (2020); Izacard et al. (2022). Structural retrieval methods retrieve structural knowledge such as neighboring entities Edge et al. (2024); Jiang et al. (2023); Wang et al. (2024) by applying graph traversal and graph machine learning Tian et al. (2024); Yasunaga et al. (2021). Despite the advancements in both textual and structural retrieval, they are often applied independently and fail to mutually reinforce each other. As shown by Figure 1(b), neither structural retrieval by following the logical structure of the query nor textual retrieval by conducting Top-K BM25 matching can achieve better performance on both Amazon and MAG datasets simultaneously.
@@
To effectively retrieve both textual and structural knowledge from TG-KBs, recent works Xia et al. (2024); Li et al. (2024) aggregate neighboring documents to fuse structural knowledge into textual narratives followed by textual retrieval, with Xia et al. (2024) filtering irrelevant neighbors by their relations and Li et al. (2024) weighted aggregating neighbors based on their fields. However, three challenges remain. First, rewording aggregated neighbors requires frequently invoking LLMs, resulting in prohibitive resources for long documents with exponentially growing neighbors. Second, structural signals humans use to form logical plans are completely discarded after neighbor aggregation. Third, rigid neighbor aggregation overlooks varying desires for structural and textual knowledge for different queries and TG-KBs. In Figure 1(c), even within MAG, queries answered by textual retrieval (i.e., $\mathcal{Q}^{\text{Text}}$) are different from queries answered by structural retrieval (i.e., $\mathcal{Q}^{\text{Struct}}$).
@@
To address the above three challenges, we infuse the mixture-of-expert philosophy into retrieval design and propose a Mixture of Structural-and-Textual Retrieval (MoR) in Figure 2. MoR begins with a planning module that generates planning graphs outlining query logics that preserve structural signals without rewording aggregated neighbors, overcoming the first and second challenges. Next, MoR interleaves structural traversal and textual matching in the reasoning module, enabling these two retrieval to reinforce each other. Finally, MoR devises a structure-aware reranker in the organization module that adaptively adjusts the importance of retrieved textual/structural knowledge, addressing the third challenge. Via Planning–Reasoning–Organizing, MoR intelligently retrieves structural and textual knowledge based on query logical structure. Our key contributions are:
@@ 2 Preliminary
Problem Setup: With the above notations, the investigated problem here is to retrieve entities $\mathcal{C}\subseteq\mathcal{V}$ supporting answering a given query $Q$.
@@ 3.1 Planning via Textual Graph Generation
To effectively reason over the underlying logic of queries and answer them, we propose a planning module that constructs a planning graph to capture their underlying logical structures. Unlike conventional approaches relying on rigid heuristics, e.g., shortest-path retrieval in knowledge graphs Luo et al. ; Delile et al. (2024), or step-by-step prompting of LLMs, which incurs high computational costs Sun et al. (2023); Wang et al. (2024), our method generates the entire planning graph in one shot, eliminating repeated LLM calls. More importantly, as planning graphs integrate entity restrictions encoding query-specific constraints and entity categories capturing broader logical structure, our MoR can generalize learned patterns and efficiently adapt to new queries with the same underlying logic. For example, any query with the form Papers associated with <institution> and are in the field of <field> shares the same patterns with the query in Figure 2. Below, we first formalize the planning graph and then optimize its generation.
@@ 4.3.2 Feature Ablation
Feature MAG Amazon
H@1 R@20 MRR H@1 R@20 MRR
Last Node 49.91 73.49 59.92 50.36 59.62 61.05
Full Path 58.19 75.01 67.14 52.19 59.92 62.24
@@ 4.4 Further Analysis
This section understands MoR’s behavior by examining three questions, each of which enriches our insight into MoR’s functionality and offers novel perspectives inspiring future query retrieval research.
@@
How does MoR perform on different logical structures? Figure 3 shows the average performance of MoR on each query group categorized by their logical structures, where "Others" refer to queries with undefined logical structures in Wu et al. MoR consistently outperforms structural and textual retrievers across different logical structures. Among all queries, MoR performs the worst on "P → P" queries due to the ambiguity, although well-known, uniquely caused by repeated product entities from multi-step traversal. The average-performing “Others" group underscores the utility of diverse planning strategies for the same query. Lastly, the skewed query distribution and retrieval performance across planning patterns reflect the varying nature of real-world planning needs. We hope these insights inspire research on data-centric reasoning designs and error control of planning.
@@
Figure 4: Saliency map visualization of query attention over three entities along the retrieved paths
@@
Does MoR indeed adaptively leverage the trajectory knowledge? To understand how our proposed reranker prioritizes candidates in the Top-K results, we visualize the saliency map by computing the gradient of ranking scores with respect to the textual fingerprint (TF) of three nodes along the traversed path, which quantifies their importance for answering a given query. Figure 4 illustrates this by analyzing trajectories for 100 ground-truth candidates across 100 queries on the Amazon and MAG datasets. Each dimension corresponds to a traversed node, with the final one representing the candidate itself. While the saliency score is concentrated in the last dimension for Amazon, MAG exhibits a more evenly distributed saliency pattern, where multiple nodes along the path contribute significantly to ranking score computation. This suggests that structural knowledge is more critical for answering queries in MAG, aligning with the previously observed lower performance of purely textual retrieval on MAG in Table 1. Further case studies explain why the reranker attends different nodes for different queries. In Figure 4(a), the reranker favors the last two dimensions as the rich textual restriction (i.e., "Northwest Company…" and "NFL Seattle…") aids in identifying the correct node at the corresponding reasoning step, as discussed in Section 3.2. The correct nodes, having higher similarity scores with the query, help guide the retrieval process toward the ground truth. Conversely, in Figure 4(b), since the first node ("University of Lausanne") helps narrow the search space and the last node ("frameless…") further filter candidates, both nodes have high saliency scores. Overall, our findings demonstrate that the reranker dynamically adapts its reliance on structural and textual knowledge depending on the dataset and query.
@@ 5 Related Work
Due to page limitation, a comprehensive version of the related work is attached in Appendix D.
@@ B.2 Implementation Details
Prompt for Planning Graph Generation: For planning graph generation in Section 3.1, we follow previous works Luo et al. ; Wu et al. to linearize the planning process by decomposing the planning graph into sequential reasoning paths, which can be generated by LLMs via next token prediction. Given the lack of ground-truth planning graphs for training, we prompt LLMs to synthesize these ground-truth planning graphs due to their superior reasoning capability. The prompt for generating the ground-truth planning graph is:
@@ Appendix C Additional Results
C.1 Query Pattern Analysis
@@
Appendix D Comprehensive Related Work
D.1 Retrieval-augmented Generation (RAG)

### v4

@@ Abstract
Text-rich Graph Knowledge Bases (TG-KBs) have become increasingly crucial for answering queries by providing textual and structural knowledge. However, current retrieval methods often retrieve these two types of knowledge in isolation without considering their mutual reinforcement and some hybrid methods even bypass structural retrieval entirely after neighboring aggregation. To fill in this gap, we propose a Mixture of Structural-and-Textual Retrieval (MoR) to retrieve these two types of knowledge via a Planning-Reasoning-Organizing framework. In the Planning stage, MoR generates textual planning graphs delineating the logic for answering queries. Following planning graphs, in the Reasoning stage, MoR interweaves structural traversal and textual matching to obtain candidates from TG-KBs. In the Organizing stage, MoR further ranks fetched candidates by their structural trajectory. Extensive experiments demonstrate the superiority of MoR in harmonizing structural and textual retrieval with discovered insights, including uneven retrieving performance across different query logics and the benefits of integrating structural trajectories for candidate reranking. Our code is available at https://github.com/Yoega/MoR.
@@ 1 Introduction
Since supporting knowledge in TG-KBs typically exhibits in both the textual and structural formats Kolomiyets and Moens (2011); Jin et al. (2024b), retrieval methods should be tailored to both formats effectively as shown in Figure 1(a). Textual retrieval methods retrieve textual knowledge such as indexed documents based on its similarity to the given query and can be broadly categorized into lexical methods (e.g., BM25) and semantic methods (e.g., Contriever) Karpukhin et al. (2020); Izacard et al. (2022). Structural retrieval methods retrieve structural knowledge such as neighboring entities Jiang et al. (2023); Edge et al. (2024); Wang et al. (2024) by applying graph traversal and graph machine learning Yasunaga et al. (2021); Tian et al. (2024). Despite the advancements in both textual and structural retrieval, they are often applied independently and fail to mutually reinforce each other. As shown by Figure 1(b), neither structural retrieval by following the logical structure of the query nor textual retrieval by conducting Top-K BM25 matching can achieve better performance on both Amazon and MAG datasets simultaneously.
@@
To effectively retrieve both textual and structural knowledge from TG-KBs, recent works Xia et al. (2024); Li et al. (2024) aggregate neighboring documents to fuse structural knowledge into textual narratives, followed by textual retrieval, with Xia et al. (2024) filtering irrelevant neighbors by their relations and Li et al. (2024) weighted aggregating neighbors based on their fields. However, three challenges remain. First, rewording aggregated neighbors requires frequently invoking LLMs, resulting in prohibitive resources for long documents with exponentially growing neighbors. Second, structural signals humans use to form logical plans are completely discarded after neighbor aggregation. Third, rigid neighbor aggregation overlooks varying desires for structural and textual knowledge for different queries and TG-KBs. Even within the same TG-KB, such as MAG in Figure 1(c), queries answered by textual retrieval (i.e., $\mathcal{Q}^{\text{Text}}$) are different from those by structural retrieval (i.e., $\mathcal{Q}^{\text{Struct}}$).
@@
To address the above three challenges, we infuse the mixture-of-expert philosophy into retrieval design and propose a Mixture Of Structural-and-Textual Retrieval (MoR) in Figure 2. MoR begins with a planning module that generates planning graphs to outline query logics and preserve structural signals without rewording aggregated neighbors, overcoming the first and second challenges. Next, MoR interleaves structural traversal and textual matching in the reasoning module, enabling these two retrievals to reinforce each other. Finally, MoR devises a structure-aware reranker in the organization module that adaptively adjusts the importance of retrieved textual/structural knowledge, addressing the third challenge. Via Planning–Reasoning–Organizing, MoR intelligently retrieves structural and textual knowledge based on query logical structure. Our key contributions are:
@@ 2 Preliminary
Problem Setup: With the above notations, the investigated problem is to retrieve entities $\mathcal{C}\subseteq\mathcal{V}$ answering a given query $Q$.
@@ 3.1 Planning via Textual Graph Generation
To effectively reason over the underlying logic of queries and answer them, we propose a planning module that constructs a planning graph to capture their underlying logical structures. Unlike conventional approaches relying on rigid heuristics, e.g., shortest-path retrieval in knowledge graphs Luo et al. ; Delile et al. (2024), or step-by-step prompting LLMs, which incur high computational costs Sun et al. (2023); Wang et al. (2024), our method generates the entire planning graph in one shot, eliminating repeated LLM calls. More importantly, as planning graphs integrate entity restrictions encoding query-specific constraints and entity categories capturing broader logical structure, our MoR can generalize learned patterns and efficiently adapt to new queries with the same underlying logic. For example, any query with the form Papers associated with <institution> and are in the field of <field> shares the same patterns with the query in Figure 2. Below, we first formalize the planning graph and then optimize its generation.
@@ 4.3.2 Feature Ablation
Feature MAG Amazon Prime
H@1 R@20 MRR H@1 R@20 MRR H@1 R@20 MRR
Last Node 49.91 73.49 59.92 50.36 59.62 61.05 33.52 61.95 44.15
Full Path 58.19 75.01 67.14 52.19 59.92 62.24 36.41 63.48 46.92
@@ 4.4 Further Analysis
This section understands MoR’s behavior by examining three questions, each of which enriches our insight into MoR’s functionality and offers novel perspectives inspiring future query retrieval research. More analysis can be found in Appendix C.
@@
How does MoR perform on different logical structures? Figure 3 shows the average performance of MoR on each query group categorized by their logical structures, where "Others" refer to queries with undefined logical structures in Wu et al. MoR consistently outperforms structural and textual retrievers across different logical structures. Among all queries, MoR performs the worst on "P → P" queries due to the ambiguity uniquely caused by repeated "product" entities from multi-step traversal. The average-performing “Others" group underscores the utility of diverse planning strategies for the same query. Lastly, the skewed query distribution and retrieval performance across planning patterns reflect the varying nature of real-world planning needs. We hope these insights inspire data-centric reasoning and error control of planning for heterogeneous query structures.
@@
Figure 4: Saliency map visualization of query attention over three entities along the retrieved paths.
@@
Does MoR indeed adaptively leverage the trajectory knowledge? To understand how our proposed reranker prioritizes candidates in the Top-K results, we visualize the saliency map by computing the gradient of ranking scores with respect to the textual fingerprint (TF) of three nodes along the traversed path, which quantifies their importance for answering a given query. Figure 4 illustrates this by analyzing trajectories for 100 ground-truth candidates across 100 queries on the Amazon and MAG datasets. Each dimension corresponds to a traversed node, with the final one representing the candidate itself. While the saliency score is concentrated in the last dimension for Amazon, MAG exhibits a more evenly distributed saliency pattern, where multiple nodes along the path contribute significantly to ranking score computation. This suggests that structural knowledge is more critical for answering queries in MAG, aligning with the previously observed lower performance of purely textual retrieval on MAG in Table 1. Further case studies explain why the reranker attends different nodes for different queries. In Figure 4(a), the reranker favors the last two dimensions as the rich textual restriction (i.e., "Northwest Company…" and "NFL Seattle…") aids in identifying the correct node at the corresponding reasoning step, as discussed in Section 3.2. These correct nodes with higher similarity scores with the query help guide the retrieval process toward the ground truth. Conversely, in Figure 4(b), since the first node ("University of Lausanne") helps narrow the search space and the last node ("frameless…") further filter candidates, both nodes have high saliency scores. Overall, our findings demonstrate that the reranker dynamically adapts its reliance on structural and textual knowledge depending on the dataset and query.
Figure 5: Traversal Identifier analysis in Top-20 retrieved candidates and information ratio in dataset.
Does MoR adaptively prioritize retrieval methods for different datasets? To assess the degree to which MoR adaptively leverages structural and textual retrieval across datasets, we calculate the following three ratios: Structure/All, Text/All: Fraction of structurally or textually retrieved candidates within the Top-20 retrieved candidates; Structure/Answer, Text/Answer: Fraction of ground-truth candidates that are structurally or textually retrieved among the Top-20; Structure (Information), Text (Information): Ratio of word counts of sampled relation/property requirements used in query construction.
The results are shown in Figure 5. On the MAG dataset, a majority of the Top-20 retrieved candidates (67%) and over 95% of the correct answers originate from structural retrieval. This underscores the critical role of structural signals in this domain. The finding is consistent with the inherently rich structural nature of MAG, evidenced by the higher proportion of Structure (Information) compared to Texture (Information), and is further supported by the substantial performance gains observed when structural features are incorporated, as demonstrated in Table 2 and Table 3. For the Amazon dataset, textual matching accounts for 52% of the Top-20 candidates and contributes more answers than structural traversal. This aligns with the rich textual content observed in Amazon, as shown in Figure 5. Overall, these results demonstrate that MoR exhibits adaptive retrieval behavior across datasets, effectively prioritizing the most informative retrieval strategy based on the underlying data characteristics. This provides strong evidence that MoR can dynamically coordinate between structural and textual retrieval sources.
4.5 Efficiency and Scalability of MoR
Method Planning Reasoning Organizing
Theoretical $\mathcal{O}(K\cdot D)$ $\mathcal{O}(Ad^{L-1})$ $\mathcal{O}(Bd^{L-1})$
Empirical (s) 0.971 0.463 0.0340
Empirical+Batch (s) 0.0134
Empirical+Parallel (s) 0.398
Table 4: Theoretical and empirical time complexity for each component during single-query retrieval. Batch and parallel optimization are applied to the Organizing and Reasoning stages, respectively.
Dataset 1-hop 2-hop 3-hop
Prime 0.059 s 0.068 s 0.072 s
Amazon 0.381 s 0.403 s 0.669 s
MAG 0.289 s 0.297 s 1.233 s
Table 5: Empirical traversal time across different hop depths for each dataset.
Since MoR consists of three stages, we conduct the complexity analysis by analyzing each of these stages theoretically and empirically. The results are shown in Table 4, where $K$ means the number of tokens in the generated planning graph, $D$ is the model dimension following the GPT-style decoding and key-value caching, $d$ denotes the average degree of node at each step, $L$ means the number of steps/layers in one path, and both $A$ and $B$ are constant. Due to the limitations of LLMs, it is hard to improve the efficiency of Planning. As the traversal is independent, we have implemented the parallel version to speed up Reasoning. For the Organizing stage, we can easily use batch to compute the ranking scores for multiple candidates and queries simultaneously, improving the efficiency.
Based on the above analysis, the time complexity for the whole framework is $\mathcal{O}(K\cdot D+(A+B)d^{L-1})\approx\mathcal{O}(d^{L-1})$, indicating that the complexity exponentially grows as the length of the reasoning path increases. To examine how time complexity varies with path length, we empirically analyze the scalability of MoR with respect to the depth of reasoning paths. We group test queries by the number of reasoning hops (e.g., 1-hop, 2-hop, 3-hop) and compare efficiency across these three groups on three datasets. As shown in Table 5, most queries can be processed within 1s. While deeper queries naturally require more processing time, MoR maintains significantly high efficiency due to its step-wise traversal and adaptive fallback to textual retrieval when necessary.
More detailed analysis of computational complexity and scalability to large datasets can be found in Appendix D.1 and Appendix D.2.
@@ 5 Related Work
Due to page limitation, a comprehensive version of the related work is attached in Appendix F.
@@ B.2 Implementation Details
Planning Graph Generation: In Section 3.1, we follow previous works Luo et al. ; Wu et al. to linearize the planning process by decomposing the planning graph into sequential reasoning paths, which can be generated by LLMs via next token prediction. Given the lack of ground-truth planning graphs for training, we prompt LLMs to synthesize these ground-truth planning graphs due to their superior reasoning capability. Since the nodes in planning graphs are entity categories/types, we include Entity Type List in the prompt. We also leverage in-context learning to help LLMs learn the mapping between the query and the corresponding planning graph. The detailed prompt for generating ground-truth planning graphs and parameters for fine-tuning the planning graph generator (i.e., Llama 3.2-3B) are shown in Prompt 1 and Table 8.
@@
Parameter Value Description
per_device_train_batch_size 4 Number of training samples per device.
gradient_accumulation_steps 8 Accumulates gradients over 8 steps.
num_train_epochs 100 Number of full passes on training dataset.
max_steps 1000 Maximum number of training steps; training stops once reached.
learning_rate 2e-4 Initial learning rate for the optimizer.
warmup_steps 5 Number of warm-up steps.
optim adamw_8bit 8-bit AdamW for efficient training.
lr_scheduler_type linear Linear decay of learning rate.
weight_decay 0.01 L2 regularization to prevent overfitting.
seed 3407 Random seed for reproducibility.
Table 8: Hyperparameter for planning graph generator.
@@ Appendix C Additional Results
C.1 Planning Performance
To investigate the performance of our planning graph generator, we conduct an experiment to evaluate the quality of the generated planning graphs for queries. We view the planning graph as correct if it enables the retrieval of at least one ground-truth candidate by traversal following its structure. In Table 9, most planning graphs are correct on MAG and Amazon datasets, demonstrating their effectiveness in guiding the traversal. Even for the domain-specific Prime dataset, more than 60% of planning graphs lead to correct retrieval, showing that LLMs still generate accurate plans and effectively support downstream reasoning and reranking.
Dataset MAG Amazon Prime
Accuracy (%) 88.85 70.03 60.44
Table 9: Performance of planning graph generator on three datasets.
C.2 Query Pattern Analysis
@@
Appendix D Efficiency Analysis
D.1 Computational Complexity of MoR
Since our Planning-Reasoning-Organizing framework is multi-staged, we theoretically/empirically analyze the time complexity of each component:
  • Planning: the planning component generates the textual planning graph by linearized token generation using a pre-trained language model, which is the well-established technique for most LLMs. Since our planning graph is category-based and its textual description consists of simple terms extracted directly from the query, the resulting textual graph is relatively simple and maintains a short token length. Therefore, the generation is highly efficient and the time complexity is linear in the number of tokens, i.e., $\mathcal{O}(K\cdot D)$ with $K$ as the number of planning graph tokens and $D$ as the model dimension, following GPT-style decoding and key-value caching.
  • Reasoning: the reasoning component conducts layer-wise breadth-first traversal following generated textual planning graphs, the time complexity of which grows exponentially with the layers/hops of traversal. Let $L$ be the maximum number of layers/hops across all reasoning paths, $c$ be the total number of distinct reasoning paths decomposed from each planning graph, $d$ be the average node degree, and $t$ denote the number of nodes retrieved via textual similarity at each layer. We begin our mixed traversal from a set of $s$ seed nodes and perform layerwise expansion by retrieving neighbors of nodes from the previous layer (constrained by the prescribed node category at that corresponding layer in the planning graph). From the second layer onward, we additionally incorporate the top-$t$ nodes ranked by textual similarity to the original query and these textually retrieved nodes also contribute neighbor expansion in the next layer. Assuming $z$ represents the time used for textual matching and node entity constraint checking, the total time complexity for reasoning by following $c$ reasoning paths is
    $\mathcal{O}(c(\underbrace{(((sd+t)d+t)...)d+t}_{L})z)$ (8)
    $=\mathcal{O}(csd^{L-1}z+ct\frac{d^{L-1}-1}{d-1}z)$
    $\approx\mathcal{O}(Ad^{L-1})$
    as $s,t,c,z$ are typically constants with small values compared with the exponentially growing paths and are absorbed together into $A$.
  • Organizing: For each candidate given by the retrieved path from reasoning stage, the organization module calculates the ranking score, which takes $\mathcal{O}(Bd^{L-1})$ with $d^{L-1}$ denotes the total number of reasoning paths and $B$ denotes the time for calculating ranking score with the defined neural network for each path.
Taking the above analysis together, the time complexity of the entire framework is $\mathcal{O}(KD+(A+B)d^{L-1})\approx\mathcal{O}(d^{L-1})$. Since our planning graph generation follows the typical LLM decoding mechanism, it does not introduce additional time overhead. Therefore, we do not focus on optimizing this stage. Instead, we optimize the Reasoning and Organizing stages, considering the exponentially growing number of retrieved paths:
  • Reasoning: we parallelize the traversal process, as each reasoning path is explored independently.
  • Organizing: we simultaneously evaluate a batch of reasoning paths and their candidates.
We report the theoretically analyzed and empirically verified time complexity for one query retrieval in Table 4. To account for differences across datasets and reduce the impact of stochasticity in the experiments, we perform three runs per dataset and report the average efficiency. We can see that despite the exponentially growing time complexity of reasoning and organizing, the empirical time consumption is rather low.
D.2 Scalability of Mixed Traversal to Large-scale TG-KBs
Our reasoning process conducts layer-wise breadth-first traversal following generated textual planning graphs, the time complexity of which grows exponentially with the increase of traversal depth. The above analysis shows that the time cost (i.e., $\mathcal{O}(Ad^{L-1})$) is mainly governed by the average node degree $d$ and the reasoning hop $L$. Therefore, even for large-scale TG-KBs, the traversal remains efficient since most graphs are sparse and the reasoning paths are reasonably short, both of which have been empirically verified in  Wu et al. .
Moreover, the three TG-KBs in our experiments are already large in scale while our reasoning traversal method achieves significant efficiency in addressing each query (see Table 7 and Table 10). Furthermore, thanks to the highly parallelizable nature of our traversal process, this runtime can be further reduced by leveraging multi-processing over traversing multiple reasoning paths.
Dataset Sequential (s) Parallel (s)
Amazon 0.416 0.271
MAG 0.891 0.846
Prime 0.081 0.076
Table 10: Sequential vs. parallel processing time of different datasets.
Appendix E Extending to Question Answering Task
Prior studies have shown that improved retrieval often leads to improved question-answering outcomes Mao et al. (2020); Lewis et al. (2020). To verify this positive correlation, we additionally evaluate the QA performance based on the retrieved information. Specifically, we use LLM GPT-4o as a judge to assess which retrieval method provides more helpful context for answering the question, allowing us to examine whether improvements in retrieval quality can convert to better QA performance. To mitigate hallucination, we enable LLM to respond with "I don’t know" when neither retrieved candidate supports an answer. We sample 100 queries for each of the three TG-KBs and retrieve top-5 candidates by using both our proposed MoR and the classical textual matching baseline BM25. To mitigate the impact of positional bias, we evaluate two different presenting orders for candidates retrieved by the two methods, resulting in 200 evaluations per dataset. The win ratios (i.e., the percentage of preferences) for both methods across three datasets are reported in Table 11.
      Dataset       BM25       MoR
      MAG       27.0%       64.0%
      Prime       33.0%       56.5%
      Amazon       40.0%       42.5%
Table 11: Comparison of BM25 and MoR QA performance across datasets.
Appendix F Comprehensive Related Work
F.1 Retrieval-augmented Generation (RAG)